### v2

# CADSpotting: Robust Panoptic Symbol Spotting on Large-Scale CAD Drawings

Abstract
We introduce CADSpotting, an efficient method for panoptic symbol spotting in large-scale architectural CAD drawings. Existing approaches struggle with the diversity of symbols, scale variations, and overlapping elements in CAD designs. CADSpotting overcomes these challenges by representing each primitive with densely sampled points, described by essential attributes like coordinates and colors. Building upon a unified 3D point cloud model for joint semantic, instance, and panoptic segmentation, CADSpotting learns robust feature representations. To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS). Moreover, we introduce a large-scale CAD dataset named LS-CAD to support our experiments. Each floorplan in LS-CAD has an average coverage of 1,000 m2 (versus 100m2 in the existing dataset), providing a valuable benchmark for symbol spotting research. Experimental results on FloorPlanCAD and LS-CAD datasets demonstrate that CADSpotting outperforms existing methods, showcasing its robustness and scalability for real-world CAD applications.
1 Introduction
Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details. These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process. For instance, the Shanghai Mercedes-Benz Arena utilized over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow. As CAD designs continue to grow in complexity, automating the detection and interpretation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
Despite the progress made in symbol recognition, the task of panoptic symbol spotting in CAD drawings remains under-explored. The challenges are multifaceted, arising from the vast diversity of symbol types, the need to differentiate between visually similar symbols, and the presence of overlapping elements. Additionally, variations in scale, orientation, and stylistic representation further complicate accurate identification.
Traditional methods approached symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets. More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task. Early solutions [5, 27, 39] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols. More advanced methods leverage Transformer architectures and attention mechanisms [6, 39], improving global relationship reasoning among symbols. These methods use rasterized pixel images derived from vector CAD drawings as input. Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting. Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings. Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence. Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types. As a result, they struggle to capture the full diversity of CAD graphical primitives. In addition, these methods focus on processing small to medium-scale datasets due to the absence of efficient strategies for preserving connectivity during symbol spotting across expansive drawings. By far, there is still a lack of efficient methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale. Specifically, CADSpotting does not rely on fixed graphic primitive types. Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations. Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set. We utilize a unified 3D point cloud processing model to learn robust features from the sampled points. CADSpotting then employs a streamlined Transformer decoder module for panoptic symbol spotting, effectively leveraging the detailed learned features to achieve precise segmentation. To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and efficient panoptic segmentation.
On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes. Each drawing follows the same fine-grained annotation standards as the FloorPlanCAD dataset, with an average coverage of over 1,000 square meters per floorplan. LS-CAD is the first dataset of its scale in this domain, part of which will be released to the research community under a license waiver, to stimulate further developments. Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
2 Related Work
Panoptic Image Segmentation.
Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 33, 1, 35] and instance segmentation [28, 8, 9, 13]. Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class. Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation. SAM2 [24] extends this approach to video segmentation. However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain. To address this challenge, Kirillov et al. [12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components. Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [35], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance. Following the large-model paradigm of SAM, SEEM [40] further refines segmentation capabilities using a prompt-driven design. Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings. Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
CAD Symbol Spotting.
Early efforts [22, 23, 26, 10, 36] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval. The introduction of deep learning techniques marked a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27]. However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments. To address this limitation, Fan et al. [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing. Building on this, Zheng et al. [39] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices. Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions. Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [38, 34, 14, 30]. The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques. Concurrently, Liu et al. [17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance. Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
3 Method
An overview of our CADSpotting method is provided in Fig.2. By integrating a dense point sampling strategy within a unified point cloud processing model, our approach efficiently extracts features for CAD primitives, producing a robust representation that supports various primitive types. Details are provided in Sec. 3.1. Then, a straightforward transformer decoder is applied to perform semantic, instance, and panoptic segmentation, with the corresponding loss function detailed in Sec.3.2. Additionally, we propose a novel Sliding Window Aggregation technique in Sec.3.3. Using a weighted voting strategy and Non-Maximum Suppression (NMS), this method allows our method to process expansive CAD drawings and perform large-scale panoptic segmentation.
3.1 Primitive Feature Learning
SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points. It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc. SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module. A masked-attention transformer decoder is used to complete the panoptic symbol spotting task. Although SymPoint demonstrates strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to fully capture the diverse range of CAD symbols. To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve both the efficiency and generalizability of primitive-level representations.
Dense Point Sampling.
In CAD drawings, graphical primitives are typically categorized into line types and shape types. Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses. Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction. Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater efficiency and adaptability, making it suitable for a wide range of CAD primitives. Fig.3 illustrates the process of our dense point sampling method.
Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector :
where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point. We then employ Point Transformer V3 [34] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides high efficiency and robust performance in feature extraction.
Feature Pooling.
To obtain the final feature representation for CAD graphical primitives, we apply primitive mixed pooling to transform point-wise features into primitive-wise features. Specifically, we perform pooling on densely sampled points within each primitive , aggregating the point features of each primitive into a single primitive-wise feature vector:
where denotes the total number of 2D sampled points, and represents the number of primitives. Let represent a single primitive with its corresponding dense point cloud . The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points :
This pooling method condenses a large set of dense point features into a more compact set of primitive features, effectively reducing computational complexity and storage requirements. At the same time, primitive-wise pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss. Loss calculations are also conducted at the primitive level, aligning with our final goal of performing CAD spotting tasks.
3.2 Panoptic Symbol Spotting
We utilize a simple transformer decoder to extract semantic and instance symbol spotting information from the learned primitive features. In our model, the pooled primitive features serve as key and value inputs to three consecutive layers of Transformer decoders, where the self-attention mechanism effectively captures dependencies among the input features. The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
Following OneFormer3D [14], we partition the decoder output into two parts: the first outputs represent instance symbol spotting proposals, and the subsequent outputs correspond to semantic symbol spotting. This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding. Here, is manually defined, and corresponds to the number of semantic categories.
We translate decoder outputs to semantic proposals and instance proposals. A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask. Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance. During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [32] to suppress non-maximal instance predictions.
Loss Function.
The loss function is composed of classification loss and primitive mask loss. The classification loss is defined as a multi-class cross-entropy loss. The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between each predicted instance and its ground truth counterpart.
3.3 Sliding Window Aggregation
To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales. During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals. However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach. Nonetheless, CAD instance symbol segmentation offers certain advantages over general image instance segmentation. Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied. Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
Therefore, we propose using a fixed-size window based on the training image dimensions and apply 2D sliding window inference on the large input drawings. For each sliding window, once the primitives intersects with or are encompassed by the window, we utilize the features of these primitives along with the window’s semantic and instance proposals, thereby increasing the likelihood of fully capturing potential instances. After gathering proposals across all windows, we apply distinct aggregation strategies for semantic and instance proposals.
For semantic proposals, we use a voting scheme to predict each primitive’s semantic label, treating each primitive as an independent unit. Primitives that extend across larger spatial areas may be captured by multiple windows. In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
For instance proposals, we directly apply Non-Maximum Suppression (NMS) to the instance proposals obtained from all windows. A potential instance may be partially or fully observed across multiple windows. By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window. Incomplete proposals from other windows are subsequently filtered out during the NMS process.
| Building | Area() | #Primitive | #Instance |
| Office 1 | 3,912 | 3,910 | 406 |
| Office 2 | 1,799 | 28,193 | 1,185 |
| Office 3 | 2,994 | 16,803 | 559 |
| Campus 1 | 1,949 | 12,571 | 603 |
| Hotel 1 | 1,193 | 26,709 | 335 |
4 Experiments
4.1 LS-CAD Dataset
We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings. Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand. Table.1 showcases five large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, stairs, elevators, different types of furniture, and parking spaces. We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset. A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
4.2 Experimental Settings
Implementation Details.
We use the train split of the FloorPlanCAD dataset to train our model. The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes. For dense point sampling, we set a fixed distance . Model training is performed on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs. We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05. Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to . We set loss weight as . Additionally, we set as the number of semantic labels, and both the top- value and are set to 220. We use as the step size of Sliding Window Aggregation on LS-CAD dataset.
Metrics.
Following the approach of [6], we assess the performance of our model using multiple metrics. For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU). For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP). For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). PQ, in particular, is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks. Further details on metric formulation are available in [6].
4.3 Quantitative Evaluation
| Methods | AP50 | AP75 | mAP |
| DINO [37] | 64.0 | 54.9 | 47.5 |
| SymPoint [16] | 66.3 | 55.7 | 52.8 |
| SymPoint-V2 [17] | 71.3 | 60.7 | 60.1 |
| CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
| Method | Total | Thing | Stuff | ||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| PanCADNet [5] | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
| CADTransformer [6] | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
| GAT-CADNet [39] | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
| SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
| SymPoint-V2 [17] | 90.1 | 96.3 | 93.6 | 90.8 | 96.6 | 94.0 | 80.8 | 90.9 | 88.9 |
| CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |
To demonstrate the effectiveness of our approach, we compare it against SOTA and concurrent methods on the FloorPlanCAD [5] and LS-CAD datasets. A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
Semantic Symbol Spotting.
As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [39], SymPoint [16] and the concurrent SymPoint-V2 [17]. Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% and SymPoint-V2 by 8.6% in mIoU.
Instance Symbol Spotting.
We also evaluate our method against DINO [37], SymPoint and SymPoint-V2 for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint. As illustrated in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 0.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint-V2.
Panoptic Symbol Spotting.
We then compare our method with the same SOTA and concurrent methods used in the semantic symbol spotting comparison. The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ. Our method achieves results comparable to the concurrent SymPoint-V2, with only a slight performance gap of no more than 1.2% , except for RQ in the Stuff type, where our method outperforms. Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
We also compare CADSpotting with SymPoint on our LS-CAD dataset. Since the original SymPoint cannot directly process large-scale CAD drawings, we first apply straightforward block partitioning, using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks for input into SymPoint, referred to as SymPointBP. The results in Table.7 show that our method with block partitioning (CADSpottingBP) achieves the highest performance in SQ. Our method with the SWA technique (CADSpottingSWA) achieves results comparable to SymPointBP and CADSpottingBP, with an absolute improvement of 2.5% in SQ compared to SymPointBP and a slight decrease of 0.8% compared to CADSpottingBP. However, CADSpottingSWA significantly outperforms SymPointBP in PQ and RQ, with an improvement of at least 20%, demonstrating that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
| Primitive Pooling Type | PQ | SQ | RQ |
| w/o Pooling | 74.6 | 89.3 | 83.6 |
| Max | 86.1 | 94.1 | 91.5 |
| Average | 87.6 | 94.5 | 92.7 |
| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
4.4 Qualitative Comparison
We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset. Fig.4 and Fig.5 show that our method achieves accurate and robust semantic and panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols. This demonstrates the effectiveness of our proposed dense point sampling based primitive features. Additionally, Fig.6 presents our semantic and panoptic symbol spotting results on a large-scale CAD drawing of an office building with 28,193 primitives and 1,185 instances from our LS-CAD dataset. These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
4.5 Ablation Studies
Feature Learning.
We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive. SymPoint employs Point Transformer(PTv1) [38] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance. To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [34], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting. As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ. When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
| Methods | Backbone | PQ | SQ | RQ |
| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
| Method | PQ | SQ | RQ |
| SymPoint + Block Partitioning | 33.5 | 84.6 | 39.6 |
| Ours + Block Partitioning | 50.1 | 87.9 | 56.6 |
| Ours + SWA | 54.3 | 87.1 | 61.8 |
Pooling.
We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer. As shown in Table 5, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ. Based on these results, we select mixed pooling for our CADSpotting approach.
Partitioning.
Table.7 shows that our method with the SWA technique (CADSpottingSWA) yields results comparable to our method using block partitioning (CADSpottingBP), with only a slight decrease of 0.8% in SQ. However, CADSpottingSWA surpasses CADSpottingBP in PQ and RQ by a margin, highlighting the effectiveness of the Sliding Window Aggregation technique in panoptic symbol spotting for large-scale CAD drawings.
5 Conclusion
We have introduced CADSpotting, a novel approach for panoptic symbol spotting in large-scale CAD drawings. Our method utilizes dense point sampling within a unified point cloud processing framework to effectively learn robust feature representations of CAD primitives. To address the challenges of precise segmentation in complex, large-scale floorplans, we propose a Sliding Window Aggregation technique, which combines weighted voting with Non-Maximum Suppression to achieve precise symbol detection and segmentation. Additionally, we present LS-CAD, a new large-scale dataset containing 50 annotated CAD floorplans from diverse building types, providing a valuable benchmark for the CAD symbol spotting task. Experimental results on both the LS-CAD and FloorPlanCAD datasets demonstrate the effectiveness, scalability, and robustness of our approach.
Despite these advancements, one limitation of our method is its panoptic symbol spotting performance on large-scale CAD drawings, which currently lags behind its performance on smaller-scale datasets like FloorPlanCAD. Addressing this gap is a focus of our ongoing and future work. We also plan to extend our approach to joint symbol spotting and 3D model generation across multiple building layers, moving towards the goal of fully automated 3D building reconstruction from CAD drawings.
Supplementary Material
Appendix A LS-CAD Dataset
In Figure 8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset. The top panel illustrates Office 3, as described in the main text, which spans a floor area of 2,994 and includes 16,803 primitives and 559 instances. The bottom panel depicts Hotel 1, also described in the main text, which spans a floor area of 1,193 and includes 26,709 primitives and 335 instances.
Appendix B Quantitative Evaluation
Detailed Ablation Study Results
We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods. Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes. The detailed results are shown in Table 9.
Appendix C Qualitative Evaluation
We provide visual comparisons of prediction results on different datasets. For FloorplanCAD, we compare the ground truth (GT) and predicted results using their primitive colors to intuitively evaluate prediction accuracy, as shown in Figure 9 and 10. For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Figure 11 and 12.
Appendix D Inference Time
We conducted a test and comparison of inference speeds for different methods on the FloorplanCAD test set using a single A6000 GPU. The results are presented in Table 8.
| Method | Backbone | Pooling | Speed |
| Handcrafted Features | PTv1 | 83ms | |
| Handcrafted Features | PTv3 | 60ms | |
| Ours | PTv3 | 133ms | |
| Ours | PTv3 | 90ms |
Appendix E Slding Window Implementation Details
In the matrix-NMS [32] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix. However, the number of instance proposals grows substantially with the size of the input drawing. If a dense matrix representation is used, both the time and memory complexities of this operation become prohibitively high.
To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows. Figure 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation. By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
| class | A | B | C | D | ||||||||
| PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
| single door | 91.59 | 97.08 | 94.34 | 75.67 | 90.68 | 83.44 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
| double door | 93.51 | 96.93 | 96.47 | 75.14 | 88.76 | 84.65 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
| sliding door | 96.54 | 98.78 | 97.73 | 83.55 | 50.31 | 83.55 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
| folding door | 57.67 | 87.88 | 65.62 | 39.66 | 95.83 | 41.38 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
| revolving door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| rolling door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| window | 89.60 | 97.24 | 92.14 | 75.49 | 91.87 | 82.17 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
| bay window | 48.15 | 95.26 | 50.55 | 41.92 | 88.51 | 47.37 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
| blind window | 90.57 | 98.17 | 92.25 | 77.71 | 96.14 | 80.83 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
| opening symbol | 46.64 | 76.82 | 60.71 | 37.97 | 72.60 | 52.30 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
| sofa | 88.53 | 96.29 | 91.93 | 71.82 | 82.05 | 87.53 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
| bed | 89.79 | 92.47 | 97.11 | 78.30 | 83.76 | 93.47 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
| chair | 83.58 | 94.04 | 88.88 | 70.05 | 85.54 | 81.89 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
| table | 75.39 | 88.46 | 85.22 | 53.18 | 78.36 | 67.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
| TV cabinet | 95.23 | 97.57 | 97.60 | 82.85 | 86.79 | 95.46 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
| Wardrobe | 93.76 | 97.23 | 96.43 | 84.53 | 89.58 | 94.36 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
| cabinet | 81.40 | 91.16 | 89.28 | 65.83 | 81.45 | 80.82 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
| gas stove | 97.34 | 99.17 | 98.16 | 90.08 | 96.15 | 93.69 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
| sink | 87.59 | 95.21 | 92.00 | 76.13 | 91.10 | 83.57 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
| refrigerator | 91.96 | 96.24 | 95.55 | 71.86 | 81.87 | 87.78 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
| air conditioner | 86.53 | 99.02 | 87.39 | 78.66 | 96.36 | 81.64 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
| bath | 79.30 | 93.31 | 84.98 | 63.84 | 83.21 | 76.73 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
| bath tub | 85.00 | 91.92 | 92.47 | 70.96 | 82.90 | 85.61 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
| washing machine | 89.17 | 97.74 | 91.23 | 78.29 | 90.75 | 86.27 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
| squat toilet | 95.35 | 98.25 | 97.05 | 78.70 | 94.34 | 83.42 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
| urinal | 94.32 | 98.76 | 95.51 | 80.33 | 92.08 | 87.24 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
| toilet | 95.21 | 98.08 | 97.07 | 81.39 | 91.92 | 88.55 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
| stairs | 86.47 | 94.28 | 91.72 | 74.68 | 89.68 | 83.27 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
| elevator | 96.15 | 97.53 | 98.58 | 79.08 | 94.76 | 83.44 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
| escalator | 73.73 | 88.37 | 83.44 | 67.98 | 79.15 | 85.88 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
| row chairs | 85.16 | 94.94 | 89.71 | 82.45 | 92.43 | 89.21 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
| parking spot | 89.68 | 95.20 | 94.20 | 84.34 | 92.67 | 91.01 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
| wall | 82.37 | 88.83 | 92.73 | 75.08 | 81.61 | 66.81 | 81.86 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
| curtain wall | 64.80 | 89.88 | 72.09 | 64.17 | 84.45 | 75.98 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
| railing | 75.67 | 92.22 | 82.06 | 62.87 | 85.69 | 73.37 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
| total | 88.91 | 95.65 | 92.96 | 74.59 | 89.26 | 83.56 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |
A: Dense point sampling + PTv3 + Pooling(Ours). B: Dense point sampling + PTv3. C: Handcrafted features + PTv3. D: Handcrafted features + PTv1

### v3

@@
 Abstract
-We introduce CADSpotting, an efficient method for panoptic symbol spotting in large-scale architectural CAD drawings.
-Existing approaches struggle with the diversity of symbols, scale variations, and overlapping elements in CAD designs.
-CADSpotting overcomes these challenges by representing each primitive with densely sampled points, described by essential attributes like coordinates and colors.
-Building upon a unified 3D point cloud model for joint semantic, instance, and panoptic segmentation, CADSpotting learns robust feature representations.
+We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings.
+Existing approaches struggle with symbol diversity, scale variations, and overlapping elements in CAD designs.
+CADSpotting overcomes these challenges by representing primitives through densely sampled points with attributes like coordinates and colors, using a unified 3D point cloud model for robust feature learning.
 To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS).
-Moreover, we introduce a large-scale CAD dataset named LS-CAD to support our experiments.
-Each floorplan in LS-CAD has an average coverage of 1,000 m2 (versus 100m2 in the existing dataset), providing a valuable benchmark for symbol spotting research.
-Experimental results on FloorPlanCAD and LS-CAD datasets demonstrate that CADSpotting outperforms existing methods, showcasing its robustness and scalability for real-world CAD applications.
+Moreover, we introduce LS-CAD, a new large-scale CAD dataset to support our experiments, with each floorplan covering around 1,000 m2, significantly larger than previous benchmarks.
+Experiments on FloorPlanCAD and LS-CAD datasets show that CADSpotting significantly outperforms existing methods.
+We also demonstrate its practical value through automating parametric 3D reconstruction, enabling interior modeling directly from raw CAD inputs.
 1 Introduction
@@
 These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process.
-For instance, the Shanghai Mercedes-Benz Arena utilized over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow.
-As CAD designs continue to grow in complexity, automating the detection and interpretation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
+For instance, the Shanghai Mercedes-Benz Arena utilize over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow.
+As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
 Despite the progress made in symbol recognition, the task of panoptic symbol spotting in CAD drawings remains under-explored.
@@
 Additionally, variations in scale, orientation, and stylistic representation further complicate accurate identification.
-Traditional methods approached symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets.
+Traditional methods approach symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets.
 More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task.
@@
 As a result, they struggle to capture the full diversity of CAD graphical primitives.
-In addition, these methods focus on processing small to medium-scale datasets due to the absence of efficient strategies for preserving connectivity during symbol spotting across expansive drawings.
-By far, there is still a lack of efficient methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
+In addition, these methods focus on processing small to medium-scale datasets due to the absence of effective strategies for preserving connectivity during symbol spotting across expansive drawings.
+By far, there is still a lack of feasible methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
 In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale.
@@
 We utilize a unified 3D point cloud processing model to learn robust features from the sampled points.
-CADSpotting then employs a streamlined Transformer decoder module for panoptic symbol spotting, effectively leveraging the detailed learned features to achieve precise segmentation.
-To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and efficient panoptic segmentation.
+CADSpotting then employs a streamlined Transformer decoder for panoptic symbol spotting, effectively leveraging the learned features to achieve precise segmentation.
+To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and effective panoptic segmentation.
 On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes.
@@
 Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
+We also demonstrate the effectiveness of CADSpotting by automating the generation of 3D interior models through parametric modeling.
+Using precise instance-level data obtained from CADSpotting, we compute spatial parameters for architectural elements, including wall contours, door and window positions, door orientations, and pivot points.
+With these spatial parameters clearly defined, we leverage parametric reconstruction techniques within Blender to efficiently build 3D interior models directly from raw CAD data.
+This approach significantly streamlines the modeling process, enhancing productivity and accuracy.
+To the best of our knowledge, our proposed pipeline is the first comprehensive solution that seamlessly integrates CAD segmentation with automated 3D interior model reconstruction.
 2 Related Work
@@
 Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [36], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance.
-Following the large-model paradigm of SAM, SEEM [40] further refines segmentation capabilities using a prompt-driven design.
+Following the large-model paradigm of SAM, SEEM [41] further refines segmentation using a prompt-driven design.
 Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings.
@@
 Early efforts [22, 23, 26, 10, 37] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval.
-The introduction of deep learning techniques marked a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27].
+The introduction of deep learning techniques marks a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27].
 However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments.
@@
 [5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing.
+Countable object instances (e.g., windows, doors) can be clearly identified and individually counted whereas uncountable stuff (e.g., wall, railing) refers to elements that cannot be individually distinguished.
 Building on this, Zheng et al.
@@
 3 Method
-An overview of our CADSpotting method is provided in Fig.2.
-By integrating a dense point sampling strategy within a unified point cloud processing model, our approach efficiently extracts features for CAD primitives, producing a robust representation that supports various primitive types.
-Details are provided in Sec.
-3.1.
-Then, a straightforward transformer decoder is applied to perform semantic, instance, and panoptic segmentation, with the corresponding loss function detailed in Sec.3.2.
-Additionally, we propose a novel Sliding Window Aggregation technique in Sec.3.3.
-Using a weighted voting strategy and Non-Maximum Suppression (NMS), this method allows our method to process expansive CAD drawings and perform large-scale panoptic segmentation.
+An overview of our CADSpotting method is illustrated in Fig.2.
+By incorporating a dense point sampling strategy within a unified point cloud processing model, our method effectively extracts features for CAD primitives, yielding a robust representation that supports various primitive types.
+We describe each module in detail in the following sections.
 3.1 Primitive Feature Learning
@@
 A masked-attention transformer decoder is used to complete the panoptic symbol spotting task.
-Although SymPoint demonstrates strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to fully capture the diverse range of CAD symbols.
-To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve both the efficiency and generalizability of primitive-level representations.
+Although SymPoint shows strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to capture the diverse range of CAD symbols.
+Real-world CAD drawings often contain various complex elements, such as Bézier curves, rather than simple predefined line types.
+Therefore, methods that rely on handcrafted features lack sufficient robustness.
+To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve generalizability of primitive-wise representations.
 Dense Point Sampling.
@@
 Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction.
-Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater efficiency and adaptability, making it suitable for a wide range of CAD primitives.
+Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater feasibility and adaptability, making it suitable for a wide range of CAD primitives.
 Fig.3 illustrates the process of our dense point sampling method.
-Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance . Each 2D point is represented by a six-dimensional vector :
-where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point.
-We then employ Point Transformer V3 [34] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides high efficiency and robust performance in feature extraction.
+Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance .
+Each 2D point is represented by a six-dimensional vector: ,where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point.
+We then employ Point Transformer V3 [35] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides robust performance in feature extraction.
 Feature Pooling.
-To obtain the final feature representation for CAD graphical primitives, we apply primitive mixed pooling to transform point-wise features into primitive-wise features.
-Specifically, we perform pooling on densely sampled points within each primitive , aggregating the point features of each primitive into a single primitive-wise feature vector:
-where denotes the total number of 2D sampled points, and represents the number of primitives.
-Let represent a single primitive with its corresponding dense point cloud .
+To obtain the final feature representation for CAD primitives, we apply Primitive Mixed Pooling (PMP) to transform point-wise features into primitive-wise features.
+Let denote our pooling operator, which aggregates the point features into primitive-wise feature vector :
+where denotes the total number of sampled points, and represents the number of primitives.
+Let represent the single primitive with its corresponding dense point cloud .
 The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points:
-This pooling method condenses a large set of dense point features into a more compact set of primitive features, effectively reducing computational complexity and storage requirements.
-At the same time, primitive-wise pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss.
-Loss calculations are also conducted at the primitive level, aligning with our final goal of performing CAD spotting tasks.
+This pooling method condenses a large set of dense point features into a more compact set of primitive-level features, effectively reducing computational complexity and storage requirements.
+At the same time, primitive-level pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss.
+Loss calculations are also conducted at the primitive-level, aligning with our final goal of performing CAD spotting tasks.
 3.2 Panoptic Symbol Spotting
@@
 The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
-Following OneFormer3D [14], we partition the decoder output into two parts: the first outputs represent instance symbol spotting proposals, and the subsequent outputs correspond to semantic symbol spotting.
+Following OneFormer3D [14], we partition the decoder output into two parts: the first represent instance proposals, and the subsequent outputs correspond to semantic symbol spotting.
+The output dimension of the + vector is , integrating both instance and semantic information.
+The first dimensions are used to generate instance masks, while the remaining dimensions determine the semantic category of each primitive.
 This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding.
-Here, is manually defined, and corresponds to the number of semantic categories.
 We translate decoder outputs to semantic proposals and instance proposals.
@@
 Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance.
+Additionally, we employ a bipartite matching strategy based on the Hungarian algorithm to find the optimal correspondence between predicted results and ground truth labels while training.
 During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [33] to suppress non-maximal instance predictions.
@@
 The classification loss is defined as a multi-class cross-entropy loss.
-The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between each predicted instance and its ground truth counterpart.
+The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between predicted instances and ground truth.
 3.3 Sliding Window Aggregation
@@
 However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach.
-Nonetheless, CAD instance symbol segmentation offers certain advantages over general image instance segmentation.
+Nonetheless, CAD instance segmentation offers certain advantages over general image instance segmentation.
 Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied.
@@
 In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
-For instance proposals, we directly apply Non-Maximum Suppression (NMS) to the instance proposals obtained from all windows.
+For instance proposals, we utilize Sparse Non-Maximum Suppression(Sparse-NMS) on the proposals generated from all windows, which increases processing efficiency (detailed analysis is provided in the supplementary material).
 A potential instance may be partially or fully observed across multiple windows.
@@
 Incomplete proposals from other windows are subsequently filtered out during the NMS process.
+3.4 Automated 3D Interior Reconstruction
+CADSpotting delivers precise panoptic segmentation results, serving as the basis for our parametric 3D interior modeling algorithm.
+Utilizing detailed instance information and primitive positions from segmentation, we apply instance-level processing techniques to compute spatial parameters for key architectural elements, specifically walls, doors, and windows.
+For doors, we determine their positions, orientations, and pivot points by analyzing the relationships between arcs and lines within each segmented instance.
+Window positions, on the other hand, are directly extracted from their corresponding instances.
+Complex wall structures require a specialized approach: we first traverse and merge endpoints of adjacent lines within wall instances, forming closed polygons.
+These polygons are then rasterized into binary images, from which final wall contours are identified using the method from [32].
+With these spatial parameters, we reconstruct a complete 3D interior model using a parametric reconstruction pipeline in Blender.
+The necessary 3D primitives for walls, doors, and windows are manually prepared in Blender, providing reusable templates for streamlined parametric interior modeling.
+This approach significantly enhances efficiency, automating the conversion of raw CAD data into structurally accurate 3D interior models within minutes through procedural generation.
+An example result from our automated 3D interior reconstruction method is illustrated in Fig.
+4, with additional details provided in the supplementary material.
 | Building | Area() | #Primitive | #Instance |
-| Office 1 | 3,912 | 3,910 | 406 |
-| Office 2 | 1,799 | 28,193 | 1,185 |
-| Office 3 | 2,994 | 16,803 | 559 |
+| Office 1 | 1,799 | 28,193 | 1,185 |
 | Campus 1 | 1,949 | 12,571 | 603 |
@@
 4 Experiments
-4.1 LS-CAD Dataset
+4.1 The Proposed LS-CAD Dataset
 We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings.
 Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand.
-Table.1 showcases five large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
-Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, stairs, elevators, different types of furniture, and parking spaces.
+Table.1 showcases 3 large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
+Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, elevators and parking spaces.
 We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset.
@@
 Implementation Details.
-We use the train split of the FloorPlanCAD dataset to train our model.
 The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes.
+To ensure fair comparison, we utilize the same training subset from the FloorPlanCAD dataset as used by SymPoint.
+We train our model on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs.
 For dense point sampling, we set a fixed distance .
-Model training is performed on a machine with 8 NVIDIA A100 GPUs, using a batch size of 2 per GPU for 512 epochs.
 We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05.
@@
 Additionally, we set as the number of semantic labels, and both the top- value and are set to 220.
-We use as the step size of Sliding Window Aggregation on LS-CAD dataset.
+We use as the step size of SWA on LS-CAD dataset.
 Metrics.
@@
 For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ).
-PQ, in particular, is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks.
+PQ is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks.
 Further details on metric formulation are available in [6].
@@
 | SymPoint [16] | 66.3 | 55.7 | 52.8 |
-| SymPoint-V2 [17] | 71.3 | 60.7 | 60.1 |
 | CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
@@
 | SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
-| SymPoint-V2 [17] | 90.1 | 96.3 | 93.6 | 90.8 | 96.6 | 94.0 | 80.8 | 90.9 | 88.9 |
 | CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |
-To demonstrate the effectiveness of our approach, we compare it against SOTA and concurrent methods on the FloorPlanCAD [5] and LS-CAD datasets.
+To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD [5] and LS-CAD datasets.
 A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
 Semantic Symbol Spotting.
-As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [39], SymPoint [16] and the concurrent SymPoint-V2 [17].
-Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% and SymPoint-V2 by 8.6% in mIoU.
+As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [40], and SymPoint [16].
+Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% in mIoU.
 Instance Symbol Spotting.
-We also evaluate our method against DINO [37], SymPoint and SymPoint-V2 for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint.
-As illustrated in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 0.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint-V2.
+We also evaluate our method against DINO [38] and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint.
+As shown in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 5.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint.
 Panoptic Symbol Spotting.
-We then compare our method with the same SOTA and concurrent methods used in the semantic symbol spotting comparison.
+We then compare our method with the same methods used in the semantic symbol spotting comparison.
 The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ.
-Our method achieves results comparable to the concurrent SymPoint-V2, with only a slight performance gap of no more than 1.2% , except for RQ in the Stuff type, where our method outperforms.
 Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
-We also compare CADSpotting with SymPoint on our LS-CAD dataset.
-Since the original SymPoint cannot directly process large-scale CAD drawings, we first apply straightforward block partitioning, using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks for input into SymPoint, referred to as SymPointBP.
-The results in Table.7 show that our method with block partitioning (CADSpottingBP) achieves the highest performance in SQ.
-Our method with the SWA technique (CADSpottingSWA) achieves results comparable to SymPointBP and CADSpottingBP, with an absolute improvement of 2.5% in SQ compared to SymPointBP and a slight decrease of 0.8% compared to CADSpottingBP.
-However, CADSpottingSWA significantly outperforms SymPointBP in PQ and RQ, with an improvement of at least 20%, demonstrating that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
-| Primitive Pooling Type | PQ | SQ | RQ |
-| w/o Pooling | 74.6 | 89.3 | 83.6 |
-| Max | 86.1 | 94.1 | 91.5 |
-| Average | 87.6 | 94.5 | 92.7 |
-| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
+We also compare the effectiveness of Sliding Window Aggregation (SWA) and Block Partitioning (BP) methods when applied to the CADSpotting model on the LS-CAD dataset.
+We apply BP method using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks.
+When employing the BP method, we ensure that each primitive is uniquely represented by retaining only those whose starting points fall within the block.
+The results in Table 5 indicate that, compared to the BP method, SWA achieves the highest performance on PQ, with an absolute improvement of 10.4%, although there is a slight decrease of 1.0% on SQ.
+The result shows that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
 4.4 Qualitative Comparison
 We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset.
-Fig.4 and Fig.5 show that our method achieves accurate and robust semantic and panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols.
+Fig.5 shows that our method achieves accurate and robust panoptic symbol spotting, even in challenging scenarios involving walls, complex or overlapping symbols, and uncommon furniture symbols.
 This demonstrates the effectiveness of our proposed dense point sampling based primitive features.
@@
 These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
+| Method | PQ | SQ | RQ |
+| Ours + BP | 42.8 | 94.8 | 45.2 |
+| Ours + SWA | 53.2 | 93.8 | 56.6 |
+To evaluate model generalizability and demonstrate the utility of LS-CAD dataset, we also train models on a combined dataset consisting of FloorPlanCAD and LS-CAD.
+Comprehensive experimental results and analyses are provided in the supplementary material.
+| Methods | Backbone | PQ | SQ | RQ |
+| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
+| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
+| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
+| Method | PQ | SQ | RQ |
+| w/o Color Feature | 85.8 | 92.5 | 92.7 |
+| Color Feature | 88.9 | 95.6 | 93.0 |
 4.5 Ablation Studies
+In this section, we present ablation studies examining the effects of feature learning and RGB information.
+Additional ablation studies, which explore various pooling strategies and evaluate performance across all classes, are provided in the supplementary material.
 Feature Learning.
@@
 When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
-| Methods | Backbone | PQ | SQ | RQ |
-| Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
-| Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
-| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
-| Method | PQ | SQ | RQ |
-| SymPoint + Block Partitioning | 33.5 | 84.6 | 39.6 |
-| Ours + Block Partitioning | 50.1 | 87.9 | 56.6 |
-| Ours + SWA | 54.3 | 87.1 | 61.8 |
-Pooling.
-We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer.
-As shown in Table 5, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ.
-Based on these results, we select mixed pooling for our CADSpotting approach.
-Partitioning.
-Table.7 shows that our method with the SWA technique (CADSpottingSWA) yields results comparable to our method using block partitioning (CADSpottingBP), with only a slight decrease of 0.8% in SQ.
-However, CADSpottingSWA surpasses CADSpottingBP in PQ and RQ by a margin, highlighting the effectiveness of the Sliding Window Aggregation technique in panoptic symbol spotting for large-scale CAD drawings.
+RGB Color Information.
+We compare the performance of CADSpotting through the inclusion versus exclusion of color information in the FloorPlanCAD dataset.
+Table.7 shows that adding RGB information improves the PQ metric by 3.1%.
 5 Conclusion
-We have introduced CADSpotting, a novel approach for panoptic symbol spotting in large-scale CAD drawings.
-Our method utilizes dense point sampling within a unified point cloud processing framework to effectively learn robust feature representations of CAD primitives.
-To address the challenges of precise segmentation in complex, large-scale floorplans, we propose a Sliding Window Aggregation technique, which combines weighted voting with Non-Maximum Suppression to achieve precise symbol detection and segmentation.
-Additionally, we present LS-CAD, a new large-scale dataset containing 50 annotated CAD floorplans from diverse building types, providing a valuable benchmark for the CAD symbol spotting task.
-Experimental results on both the LS-CAD and FloorPlanCAD datasets demonstrate the effectiveness, scalability, and robustness of our approach.
-Despite these advancements, one limitation of our method is its panoptic symbol spotting performance on large-scale CAD drawings, which currently lags behind its performance on smaller-scale datasets like FloorPlanCAD.
-Addressing this gap is a focus of our ongoing and future work.
-We also plan to extend our approach to joint symbol spotting and 3D model generation across multiple building layers, moving towards the goal of fully automated 3D building reconstruction from CAD drawings.
+We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings.
+Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation.
+We also present LS-CAD, a dataset of 50 annotated floorplans from diverse building types.
+Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
+Our approach currently faces limitations due to the dataset’s primary emphasis on residential and commercial buildings, which restricts its applicability to industrial and cultural contexts.
+Additionally,it may struggle to capture significant stylistic variations among CAD symbols.
+Addressing these challenges through the integration of in-context learning may help enhance annotation efficiency and extend applicability.
+This is left as our future work.
 Supplementary Material
-Appendix A LS-CAD Dataset
-In Figure 8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset.
-The top panel illustrates Office 3, as described in the main text, which spans a floor area of 2,994 and includes 16,803 primitives and 559 instances.
-The bottom panel depicts Hotel 1, also described in the main text, which spans a floor area of 1,193 and includes 26,709 primitives and 335 instances.
-Appendix B Quantitative Evaluation
-Detailed Ablation Study Results
+| Method | Training Dataset | Test on FloorPlanCAD | Test on Cross-dataset | ||||
+| PQ | SQ | RQ | PQ | SQ | RQ | ||
+| SymPoint [16] | FloorPlanCAD | 83.3 | 91.4 | 91.1 | 33.2 | 83.0 | 39.9 |
+| SymPoint | Cross-dataset | 79.0 | 89.2 | 88.6 | 57.0 | 84.0 | 67.8 |
+| CADSpotting | FloorPlanCAD | 88.9 | 95.6 | 93.0 | 42.8 | 94.8 | 45.2 |
+| CADSpotting | Cross-dataset | 87.3 | 95.4 | 91.4 | 69.9 | 91.6 | 76.3 |
+| CADSpotting(SWA) | FloorPlanCAD | / | 53.2 | 93.8 | 56.6 | ||
+| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 84.9 | 88.9 | ||
+6 More Details on the LS-CAD Dataset
+In Fig.8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset.
+The top panel illustrates Office 3, which spans a floor area of 2,994, including 16,803 primitives and 559 instances.
+The bottom panel depicts Hotel 1, which spans a floor area of 1,193, consisting of 26,709 primitives and 335 instances.
+To demonstrate the generalization capability of our model and highlight the value of the LS-CAD dataset, we implement a cross-dataset joint training paradigm.
+We randomly partition complete large-scale CAD drawings into three subsets: 80% for training, 10% for validation, and 10% for testing.
+The division of the LS-CAD dataset strictly adheres to the standardized data splitting protocol established in FloorPlanCAD.
+The training and validation sets from both datasets are subsequently integrated to construct a cross-dataset collaborative training environment.
+As shown in Table.8, the cross-dataset joint training strategy significantly improves model performanc on the LS-CAD test set, although with slight performance degradation on the original FloorPlanCAD test set.
+This demonstrates the necessity of the proposed LS-CAD dataset for provding more diverse CAD representations.
+Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric.
+After cross-dataset joint training, the PQ score for CADSpotting decreases by only 1.2% on FloorPlanCAD, whereas it drops by 4.3% on SymPoint.
+This indicates that our model exhibits stronger generalization capabilities.
+7 Quantitative Evaluation
+Ablation Study on Feature Learning.
 We present a comprehensive evaluation of panoptic quality (PQ), segmentation quality (SQ), and recognition quality (RQ) for each class, comparing different feature learning methods.
 Our final approach achieves superior performance across the majority of classes, particularly excelling in commonly used categories such as door classes, window classes, and wall classes.
-The detailed results are shown in Table 9.
-Appendix C Qualitative Evaluation
-We provide visual comparisons of prediction results on different datasets.
-For FloorplanCAD, we compare the ground truth (GT) and predicted results using their primitive colors to intuitively evaluate prediction accuracy, as shown in Figure 9 and 10.
-For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Figure 11 and 12.
-Appendix D Inference Time
-We conducted a test and comparison of inference speeds for different methods on the FloorplanCAD test set using a single A6000 GPU. The results are presented in Table 8.
-| Method | Backbone | Pooling | Speed |
-| Handcrafted Features | PTv1 | 83ms | |
-| Handcrafted Features | PTv3 | 60ms | |
-| Ours | PTv3 | 133ms | |
-| Ours | PTv3 | 90ms |
-Appendix E Slding Window Implementation Details
+The detailed results are shown in Table.10.
+Ablation Study on Pooling.
+We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer.
+As shown in Table.9, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ.
+Based on these results, we select mixed pooling for our CADSpotting approach.
+8 Qualitative Evaluation
+We provide comparisons of prediction results on different datasets.
+For FloorPlanCAD, we compare the ground truth and prediction using their primitive colors to intuitively evaluate prediction accuracy, as shown in Fig.9, Fig.10 and Fig.13.
+For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Fig.11 and Fig.12.
+9 Sliding Window Implementation Details
+The Algorithm.1 details the implementation process of our SWA.
 In the matrix-NMS [33] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix.
@@
 To mitigate this issue, we observe that only a small fraction of proposal pairs exhibit nonzero overlap, as most proposals belong to different inference windows.
-Figure 7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation.
+Fig.
+7 visualizes the sparsity structure of the IoU matrix during sliding window aggregation.
 By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
+10 Automated 3D Interior Reconstruction
+Based on the spotting results generated by CADSpotting, we present more 3D reconstruction renderings in Fig.14.
+By systematically optimizing the parameters of predefined instance components using instance segmentation data and primitive spatial coordinates, we achieve accurate 3D model reconstruction through component-level geometric adaptation
+For wall polygon extraction, we employ an innovative SVG-PNG conversion strategy to differentiate floor areas from walls: 1) Merge adjacent endpoints and remove duplicate segments to simplify SVG paths.
+2) Convert vectors to 8K resolution binarized PNG while maintaining coordinate alignment.
+3) Identify the largest connected component as the floor through connected component analysis, with secondary components as candidate walls.
+4) Re-vectorize boundaries through coordinate mapping to generate closed wall polygons.
+The proposed method exhibits notable robustness against recognition inaccuracies – incomplete walls are automatically reconciled during stage 1, while small noise particles are filtered in stage 3 using area thresholds.
+For doors and windows, we employ category-specific parameterization strategies.
+Doors are systematically categorized into four distinct subtypes (single/double/sliding/folding) based on semantic labels, with positions, orientations, and pivot points determined through linear-arc spatial relationships.
+Windows are grouped via union-find algorithms, calculating average lines of co-grouped segments as representative lines, then selecting the segment closest to the representative line within each group as the window instance position.
+The final reconstruction leverages Blender’s geometry nodes for procedural modeling: 1) Extrude wall polygons to 3D volumes using floorplan height attributes.
+2) Instance-place door/window assemblies at infered positions.
+3) Apply floor based on origin CAD drawing.
+This integrated approach achieves automated 3D reconstruction with exceptional efficiency, transforming raw CAD data into structurally complete 3D interior models within minutes through streamlined procedural generation.
+| Primitive Pooling Type | PQ | SQ | RQ |
+| w/o Pooling | 74.6 | 89.3 | 83.6 |
+| Max | 86.1 | 94.1 | 91.5 |
+| Average | 87.6 | 94.5 | 92.7 |
+| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
 | class | A | B | C | D | ||||||||

### v4

@@
 We introduce CADSpotting, an effective method for panoptic symbol spotting in large-scale architectural CAD drawings.
-Existing approaches struggle with symbol diversity, scale variations, and overlapping elements in CAD designs.
-CADSpotting overcomes these challenges by representing primitives through densely sampled points with attributes like coordinates and colors, using a unified 3D point cloud model for robust feature learning.
-To enable accurate segmentation in large, complex drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining weighted voting and Non-Maximum Suppression (NMS).
-Moreover, we introduce LS-CAD, a new large-scale CAD dataset to support our experiments, with each floorplan covering around 1,000 m2, significantly larger than previous benchmarks.
-Experiments on FloorPlanCAD and LS-CAD datasets show that CADSpotting significantly outperforms existing methods.
-We also demonstrate its practical value through automating parametric 3D reconstruction, enabling interior modeling directly from raw CAD inputs.
-1 Introduction
+Existing approaches often struggle with symbol diversity, scale variations, and overlapping elements in CAD designs, and typically rely on additional features (e.g., primitive types or graphical layers) to improve performance.
+CADSpotting overcomes these challenges by representing primitives through densely sampled points with only coordinate attributes, using a unified 3D point cloud model for robust feature learning.
+To enable accurate segmentation in large drawings, we further propose a novel Sliding Window Aggregation (SWA) technique that combines weighted voting and Non-Maximum Suppression (NMS).
+Moreover, we introduce LS-CAD, a new large-scale dataset comprising 45 finely annotated floorplans, each covering approximately 1,000 , significantly larger than prior benchmarks.
+LS-CAD will be publicly released to support future research.
+Experiments on FloorPlanCAD and LS-CAD demonstrate that CADSpotting significantly outperforms existing methods.
+We also showcase its practical value by enabling automated parametric 3D interior reconstruction directly from raw CAD inputs.
+Introduction
 Architectural Computer-Aided Design (CAD) drawings, especially those used for interior design, serve as detailed digital representations that convey essential information about a building’s structure, layout, and intricate details.
 These drawings include critical elements such as furniture placement, electrical layouts, and spatial organization, ensuring consistency and precision throughout the construction process.
-For instance, the Shanghai Mercedes-Benz Arena utilize over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow.
-As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling [20].
+For instance, the Shanghai Mercedes-Benz Arena utilizes over 3,000 CAD drawings to guide its structural design, optimize spatial arrangements, and streamline the construction workflow.
+As CAD designs continue to grow in complexity, automating the detection and interpolation of symbols within these floorplans becomes increasingly important in various downstream tasks, such as code compliance checking and 3D interior modeling (MarsBIM 2024).
 Despite the progress made in symbol recognition, the task of panoptic symbol spotting in CAD drawings remains under-explored.
@@
 Additionally, variations in scale, orientation, and stylistic representation further complicate accurate identification.
-Traditional methods approach symbol recognition in a query-by-example manner [29], but these techniques struggle to handle the vast diversity of graphical symbols in real-world datasets.
-More recent learning-based approaches [25, 7] have shown significant progress in addressing the symbol spotting task.
-Early solutions [5, 27, 40] typically leverage convolutional neural networks (CNNs) or graph-based models to spot symbols.
-More advanced methods leverage Transformer architectures and attention mechanisms [6, 40], improving global relationship reasoning among symbols.
-These methods use rasterized pixel images derived from vector CAD drawings as input.
-Such conversion, however, introduces discrepancies between the original vector graphics and the rasterized images, resulting in errors due to the loss of precise geometric information.
-The representative work SymPoint [16] treats CAD drawings as sets of primary points, leveraging techniques from point cloud analysis for effective CAD symbol spotting.
-Such a representation simplifies model complexity while achieving strong performance on panoptic symbol spotting tasks in CAD drawings.
-Building on this foundation, SymPoint-V2 [17] introduces Layer feature-enhanced encoding to incorporate graphical layer information and a position-guided Training method to enhance model learning and accelerate convergence.
-Yet both approaches rely on manually defined primitive feature representations, i.e., four fixed primitive types.
-As a result, they struggle to capture the full diversity of CAD graphical primitives.
-In addition, these methods focus on processing small to medium-scale datasets due to the absence of effective strategies for preserving connectivity during symbol spotting across expansive drawings.
-By far, there is still a lack of feasible methods to address the panoptic symbol spotting task in large-scale CAD drawings, as dense symbol clutter and significant variations in scale present substantial challenges for accurate segmentation and symbol recognition.
-In this paper, we present CADSpotting, a simple yet highly effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale.
-Specifically, CADSpotting does not rely on fixed graphic primitive types.
-Instead, it densely samples points along CAD graphic primitives to construct comprehensive point cloud representations.
-Each point is represented by basic attributes, such as coordinates and color, forming a lightweight but expressive feature set.
-We utilize a unified 3D point cloud processing model to learn robust features from the sampled points.
-CADSpotting then employs a streamlined Transformer decoder for panoptic symbol spotting, effectively leveraging the learned features to achieve precise segmentation.
-To handle large-scale, real-world CAD drawings, we further propose a novel Sliding Window Aggregation (SWA) technique, combining a weighted voting strategy with Non-Maximum Suppression (NMS) to enable accurate and effective panoptic segmentation.
-On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 50 annotated floorplans from a variety of building types, including campuses and office complexes.
-Each drawing follows the same fine-grained annotation standards as the FloorPlanCAD dataset, with an average coverage of over 1,000 square meters per floorplan.
-LS-CAD is the first dataset of its scale in this domain, part of which will be released to the research community under a license waiver, to stimulate further developments.
-Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods by a margin, demonstrating its robustness and effectiveness in handling the complexities of large-scale, real-world CAD drawings.
-We also demonstrate the effectiveness of CADSpotting by automating the generation of 3D interior models through parametric modeling.
-Using precise instance-level data obtained from CADSpotting, we compute spatial parameters for architectural elements, including wall contours, door and window positions, door orientations, and pivot points.
-With these spatial parameters clearly defined, we leverage parametric reconstruction techniques within Blender to efficiently build 3D interior models directly from raw CAD data.
-This approach significantly streamlines the modeling process, enhancing productivity and accuracy.
-To the best of our knowledge, our proposed pipeline is the first comprehensive solution that seamlessly integrates CAD segmentation with automated 3D interior model reconstruction.
-2 Related Work
+Traditional query-by-example based methods (Rusiñol, Lladós, and Sánchez 2010) struggle with the diverse and complex graphical symbols in real-world CAD drawings.
+Learning-based approaches (Ren et al.
+2016; Farhadi and Redmon 2018) have advanced the field, with early methods using convolutional neural networks (CNNs) or graph-based models (Fan et al.
+2021; Rezvanifar, Cote, and Albu 2020; Zheng et al.
+2022), and later works incorporating Transformers and attention mechanisms (Fan et al.
+2022; Zheng et al.
+2022) for improved global reasoning.
+However, these approaches typically operate on rasterized images derived from vector CAD data, leading to loss of geometric fidelity and recognition errors.
+SymPoint (LIU et al.
+2024) avoids rasterization by treating CAD drawings as point sets and applying point cloud techniques for effective symbol spotting.
+SymPoint-V2 (Liu et al.
+2024) builds on this by introducing layer-aware encoding and position-guided training.
+While effective, both methods rely on predefined primitive types, and SymPoint-V2 additionally depends on graphical layer information, limiting their ability to generalize across diverse real CAD data.
+Moreover, they focus on small to medium-scale drawings and lack mechanisms for preserving spatial consistency in large-scale drawings.
+As a result, panoptic symbol spotting in expansive CAD drawings remains challenging due to dense clutter, scale variation, and segmentation complexity.
+In this paper, we present CADSpotting, a simple yet effective panoptic symbol spotting method tailored for handling CAD drawings at a much larger scale.
+Unlike prior approaches, CADSpotting does not rely on fixed graphic primitive types or additional layer information.
+Instead, it densely samples points along CAD graphic primitives to construct a compact yet expressive point cloud representation, where each point is represented solely by its coordinate attributes.
+A unified 3D point cloud model is employed to learn robust features from these sampled points, followed by a streamlined Transformer decoder that performs panoptic symbol spotting.
+To handle large-scale, real-world CAD drawings, we introduce a novel Sliding Window Aggregation (SWA) technique that combines weighted voting and Non-Maximum Suppression (NMS) to enable accurate and scalable panoptic segmentation across windowed regions.
+On the data front, we further introduce LS-CAD, a new large-scale CAD dataset containing 45 finely annotated floorplans from diverse building types, including campuses and office complexes.
+Each floorplan spans over 1,000 on average and follows the same fine-grained annotation standards as FloorPlanCAD.
+LS-CAD is the first dataset of its scale in this domain, and it will be released under a license waiver to support future research.
+Experimental results on both FloorPlanCAD and LS-CAD demonstrate that CADSpotting outperforms the state-of-the-art methods, demonstrating its robustness and effectiveness in real-world scenarios.
+Finally, we demonstrate CADSpotting’s practical value by automating 3D interior reconstruction.
+Leveraging instance-level outputs, we extract spatial parameters such as wall contours, door/window positions, orientations, and pivot points.
+These parameters are then used to drive parametric modeling in Blender, enabling efficient and accurate 3D interior generation directly from raw CAD inputs.
+To our knowledge, this constitutes the first end-to-end pipeline that tightly integrates CAD symbol spotting with parametric 3D reconstruction (Fig.1).
+Related Work
 Panoptic Image Segmentation.
-Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation [18, 2, 34, 1, 36] and instance segmentation [28, 8, 9, 13].
-Semantic segmentation assigns a class label to every pixel in an image, while instance segmentation distinguishes individual objects within the same class.
-Recently, SAM [13] introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation.
-SAM2 [24] extends this approach to video segmentation.
-However, semantic and instance segmentation methods struggle to accurately represent uncountable background elements, such as the sky or terrain.
-To address this challenge, Kirillov et al.
-[12] propose Panoptic Segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components.
-Early approaches [11, 31, 15, 3] primarily employ CNN-bases architectures, while recent advancements, such as Mask2Former [4], SegFormer [36], and OneFormer [9], leverage Transformer models, incorporating a mask attention mechanism to boost segmentation performance.
-Following the large-model paradigm of SAM, SEEM [41] further refines segmentation using a prompt-driven design.
-Despite these advances, most existing segmentation models remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings.
-Consequently, they often fail to capture critical vector-based information, such as overlapping relationships and precise geometric attributes.
+Image segmentation is a fundamental task in computer vision, traditionally divided into two main categories: semantic segmentation (Long, Shelhamer, and Darrell 2015; Chen et al.
+2017; Wu et al.
+2019; Chen 2017; Xie et al.
+2021) and instance segmentation (Ronneberger, Fischer, and Brox 2015; He et al.
+2017; Jain et al.
+2023; Kirillov et al.
+2023).
+While semantic segmentation labels each pixel by class, instance segmentation further distinguishes between object instances of the same class.
+Recently, SAM (Kirillov et al.
+2023) introduces a large model trained on over one billion masks, leveraging prompt-based interactions for enhanced interactive segmentation.
+SAM2 (Ravi et al.
+2024) extends this approach to video segmentation.
+However, these methods often struggle to represent uncountable background elements (e.g., sky, terrain).
+To unify countable and uncountable regions, Kirillov et al.
+(Kirillov et al.
+2019) introduce panoptic segmentation, which unifies semantic and instance segmentation, effectively handling both countable and uncountable background components.
+Early approaches(Sun et al.
+2019; Li et al.
+2019; Chen, Wang, and Qiao 2020) use CNN backbones, while recent advances like Mask2Former (Cheng et al.
+2022) and OneFormer (Jain et al.
+2023) leverage Transformer architectures with mask attention mechanisms for improved accuracy.
+Large models such as SEEM (Zou et al.
+2024) further refines segmentation using a prompt-driven design.
+Despite these advances, most methods remain pixel-centric, limiting their effectiveness in processing vector graphics like CAD drawings.
+As a result, they struggle to capture critical vector-based information, including precise geometry and overlapping relationships.
 CAD Symbol Spotting.
-Early efforts [22, 23, 26, 10, 37] in symbol spotting on CAD drawings rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval.
-The introduction of deep learning techniques marks a significant shift, with models based on Faster R-CNN [25] and YOLO v3 [7] greatly enhancing the accuracy of symbol recognition tasks [27].
-However, traditional symbol spotting methods are primarily designed for countable objects and do not effectively handle uncountable symbols, which are common in CAD environments.
-To address this limitation, Fan et al.
-[5] introduce the FloorPlanCAD dataset and CNN-GCN method, enabling the recognition of both countable and uncountable symbols for more comprehensive CAD parsing.
-Countable object instances (e.g., windows, doors) can be clearly identified and individually counted whereas uncountable stuff (e.g., wall, railing) refers to elements that cannot be individually distinguished.
-Building on this, Zheng et al.
-[40] propose GAT-CADNet, a model that represents CAD drawings as graph structures and performs segmentation by predicting adjacency matrices.
-Further advancements include CADTransformer [6], which utilizes the Vision Transformer to extract features from scalar maps, leading to improved predictions.
-Meanwhile, significant progress has been made in the field of 3D point cloud segmentation [39, 35, 14, 30].
-The seminal work SymPoint [16] proposes a novel method that converts symbols into primary point representations, thereby enhancing feature extraction through point cloud analysis and upsampling techniques.
-Concurrently, Liu et al.
-[17] extend SymPoint by incorporating LFE and PGT modules, further improving feature representation and model performance.
-Despite these advances, the manually defined feature representation is often constrained by a limited set of graphical primitive types, which hinders their ability to capture the full diversity of CAD symbols.
-3 Method
+Early CAD symbol spotting methods (Nguyen, Tabbone, and Terrades 2008; Nguyen, Tabbone, and Boucher 2009; Rezvanifar, Cote, and Branzan Albu 2019; Jiang et al.
+2021; Yang et al.
+2023) rely on handcrafted feature descriptors, such as shape and structure, coupled with techniques like sliding window searches or graph matching for symbol retrieval.
+The introduction of deep learning significantly improved performance, with models based on Faster R-CNN (Ren et al.
+2016) and YOLO (Farhadi and Redmon 2018) achieving higher accuracy (Rezvanifar, Cote, and Albu 2020), though they primarily focus on countable object instances.
+To address uncountable elements, Fan et al.(Fan et al.
+2021) introduce the FloorPlanCAD dataset and a CNN-GCN framework for holistic parsing.
+Zheng et al.(Zheng et al.
+2022) further model CAD drawings as graphs, using adjacency prediction for segmentation.
+CADTransformer (Fan et al.
+2022) adopt Vision Transformers to extract features from scalar maps.
+Inspired by advancements in 3D point cloud segmentation (Zhao et al.
+2021; Wu et al.
+2024; Kolodiazhnyi et al.
+2024; Schult et al.
+2023), SymPoint (LIU et al.
+2024) reformulates CAD drawings as primary point sets and applies point-based learning for effective symbol spotting.
+SymPoint-V2 (Liu et al.
+2024) extends this with layer-aware encoding and position-guided training.
+Despite their effectiveness, both methods rely on a fixed set of primitive types, and SymPoint-V2 assumes the availability of reliable layer information.
+These assumptions often do not hold in real-world CAD data, limiting their ability to capture the full diversity and complexity of CAD symbols.
+Method
 An overview of our CADSpotting method is illustrated in Fig.2.
@@
 We describe each module in detail in the following sections.
-3.1 Primitive Feature Learning
-SymPoint [16] introduces a new concept that represents CAD primitives as a set of 2D points.
-It constructs a handcrafted feature for each graphical primitive, capturing details such as type (e.g., line, arc, circle, ellipse), length, etc.
-SymPoint then applies point cloud analysis to learn primitive representations, further enhancing connectivity through an Attention with Connection Module.
-A masked-attention transformer decoder is used to complete the panoptic symbol spotting task.
-Although SymPoint shows strong performance in CAD symbol spotting, its feature representation is manually defined and limited to four fixed primitive types, which may constrain its capacity to capture the diverse range of CAD symbols.
-Real-world CAD drawings often contain various complex elements, such as Bézier curves, rather than simple predefined line types.
-Therefore, methods that rely on handcrafted features lack sufficient robustness.
-To address these constraints, we propose a novel graphical primitive feature representation for CAD drawings that incorporates dense point sampling within a unified point cloud analysis model to improve generalizability of primitive-wise representations.
-Dense Point Sampling.
-In CAD drawings, graphical primitives are typically categorized into line types and shape types.
-Line primitives encompass straight segments, curves, and complex contours such as Bézier curves, while shape primitives include basic geometric forms like rectangles, circles, and ellipses.
-Inspired by SymPoint [16], we propose a dense point sampling strategy that performs equidistant sampling on each primitive to generate dense point data for feature extraction.
-Rather than assigning fixed types to CAD graphical primitives and using handcrafted features to represent them, our dense point sampling method offers greater feasibility and adaptability, making it suitable for a wide range of CAD primitives.
-Fig.3 illustrates the process of our dense point sampling method.
-Given a CAD drawing containing primitives, we perform dense sampling to generate a 2D point cloud from these primitives at a fixed distance .
-Each 2D point is represented by a six-dimensional vector: ,where represent the 2D coordinates of the point, is set to 0, and represent the color information (red, green, blue channels) of the point.
-We then employ Point Transformer V3 [35] (PTv3) as the backbone for feature extraction, leveraging its design optimized for the unordered nature of point clouds, which provides robust performance in feature extraction.
-Feature Pooling.
-To obtain the final feature representation for CAD primitives, we apply Primitive Mixed Pooling (PMP) to transform point-wise features into primitive-wise features.
-Let denote our pooling operator, which aggregates the point features into primitive-wise feature vector :
-where denotes the total number of sampled points, and represents the number of primitives.
-Let represent the single primitive with its corresponding dense point cloud .
-The -th dimension feature of is determined by the sum of the maximum and average values of the corresponding -th dimension features across all points:
-This pooling method condenses a large set of dense point features into a more compact set of primitive-level features, effectively reducing computational complexity and storage requirements.
-At the same time, primitive-level pooling retains the essential information of the contained points, preserving global consistency in the primitive representation and minimizing information loss.
-Loss calculations are also conducted at the primitive-level, aligning with our final goal of performing CAD spotting tasks.
-3.2 Panoptic Symbol Spotting
-We utilize a simple transformer decoder to extract semantic and instance symbol spotting information from the learned primitive features.
-In our model, the pooled primitive features serve as key and value inputs to three consecutive layers of Transformer decoders, where the self-attention mechanism effectively captures dependencies among the input features.
-The queries are initialized randomly and optimized during training to compute accurate self-attention scores with other features.
-Following OneFormer3D [14], we partition the decoder output into two parts: the first represent instance proposals, and the subsequent outputs correspond to semantic symbol spotting.
-The output dimension of the + vector is , integrating both instance and semantic information.
-The first dimensions are used to generate instance masks, while the remaining dimensions determine the semantic category of each primitive.
-This configuration enables panoptic symbol spotting by integrating instance and semantic information for comprehensive scene understanding.
-We translate decoder outputs to semantic proposals and instance proposals.
-A semantic proposal includes the log-likelihood values for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask.
-Similarly, an instance proposal also contains log-likelihood values, but these indicate the likelihood of each primitive being classified as a specific instance.
-Additionally, we employ a bipartite matching strategy based on the Hungarian algorithm to find the optimal correspondence between predicted results and ground truth labels while training.
-During inference, we select the top- instance proposals of the highest confidence scores and apply matrix-NMS [33] to suppress non-maximal instance predictions.
-Loss Function.
-The loss function is composed of classification loss and primitive mask loss.
+Primitive Feature Learning
+SymPoint (LIU et al.
+2024) represents CAD primitives as sets of 2D points and constructs handcrafted features for each, encoding attributes such as type (line, arc, circle, ellipse) and length.
+It then applies point cloud analysis with an Attention with Connection Module to learn primitive representations, followed by a masked-attention Transformer decoder for panoptic symbol spotting.
+While effective, SymPoint is limited by its reliance on manually defined features and a fixed set of four primitive types, which restricts its ability to handle the full diversity of CAD symbols.
+Real-world drawings often include complex elements like Bézier curves, making handcrafted features insufficiently robust.
+To overcome these limitations, we propose a novel primitive representation based on dense point sampling and unified point cloud learning, enabling more generalizable and expressive primitive-level features.
+Dense Point Sampling
+In CAD drawings, graphical primitives are generally classified into line types (e.g., straight segments, curves, Bézier contours) and shape types (e.g., rectangles, circles, ellipses).
+Inspired by SymPoint (LIU et al.
+2024), we adopt a dense point sampling strategy that performs equidistant sampling along each primitive to generate dense point-level data for feature learning.
+Unlike methods that rely on fixed primitive types and handcrafted features, our sampling approach is more flexible and generalizable, making it suitable for a wide variety of CAD elements.
+Fig.
+3 illustrates the sampling process.
+Given a CAD drawing with primitives, we sample points at a fixed interval to construct a 2D point cloud , where each point is represented as a 3D vector with .
+We then use Point Transformer V3 (Wu et al.
+2024) (PTv3) as the backbone to extract robust features, leveraging its design optimized for unordered point cloud data.
+Feature Pooling
+To obtain the final feature representation for CAD primitives, we use Primitive Mixed Pooling (PMP) to convert point-wise features into primitive-wise features .
+where is the total number of sampled points, and is the number of primitives.
+Let denote the dense point set corresponding to the -th primitive.
+The -th dimension of the primitive-level feature is computed as the sum of the maximum and average values of the -th dimension across all points in :
+This pooling strategy compresses dense point features into a compact set of primitive-level representations, reducing computational and storage overhead.
+Meanwhile, it preserves essential geometric information, maintaining global consistency and minimizing information loss.
+Loss computation is also performed at the primitive level, aligning with the objective of CAD symbol spotting.
+Panoptic Symbol Spotting
+We utilize a lightweight Transformer decoder to extract semantic and instance-level information from the pooled primitive features.
+These features serve as keys and values for three consecutive decoder layers, where self-attention captures dependencies among primitives.
+Query embeddings are randomly initialized and optimized during training to compute accurate self-attention scores with other features.
+Following OneFormer3D (Kolodiazhnyi et al.
+2024), the decoder output is divided into two parts: the first vectors correspond to instance proposals, and the remaining vectors represent semantic symbol spotting.
+The output has shape , where the first dimensions are used to predict instance masks, and the rest indicate semantic class scores.
+This unified representation enables panoptic symbol spotting by combining both semantic and instance-level predictions.
+Each semantic proposal includes log-likelihoods for each primitive being classified into different semantic labels and can be further processed through thresholding to generate a binary mask.
+Similarly, an instance proposal also contains log-likelihood values, and these indicate the likelihood of each primitive being classified as a specific instance.
+During training, we apply bipartite matching using the Hungarian algorithm to align predictions with ground truth.
+At inference, we select the top- instance proposals with the highest confidence and apply matrix-NMS (Wang et al.
+2020) to eliminate redundant predictions.
+Loss Function The overall loss function consists of a classification loss and a primitive mask loss.
 The classification loss is defined as a multi-class cross-entropy loss.
-The primitive mask loss combines binary cross-entropy loss, , with the Dice loss [21], , to evaluate the correspondence between predicted instances and ground truth.
-3.3 Sliding Window Aggregation
-To address the challenge of panoptic symbol segmentation at large scales, a key issue is ensuring the generalization of instance symbol segmentation across varying scales.
-During inference on larger drawings than those encountered during training, the network must produce significantly more instance proposals.
-However, this process is limited by the fixed number of instance queries inherent in Mask Transformer-based methods, a constraint that also affects our approach.
-Nonetheless, CAD instance segmentation offers certain advantages over general image instance segmentation.
-Specifically, in CAD floorplans, instances of the same category maintain relatively consistent sizes across different scenes when a uniform scale is applied.
-Consequently, we can reasonably assume that the size of instances in inference will not exceed that of the same category in the training images.
-Therefore, we propose using a fixed-size window based on the training image dimensions and apply 2D sliding window inference on the large input drawings.
-For each sliding window, once the primitives intersects with or are encompassed by the window, we utilize the features of these primitives along with the window’s semantic and instance proposals, thereby increasing the likelihood of fully capturing potential instances.
-After gathering proposals across all windows, we apply distinct aggregation strategies for semantic and instance proposals.
-For semantic proposals, we use a voting scheme to predict each primitive’s semantic label, treating each primitive as an independent unit.
-Primitives that extend across larger spatial areas may be captured by multiple windows.
-In such cases, we prioritize proposals that capture the primitive as completely as possible by using weighted voting, where each window’s voting weight is determined by the proportion of the primitive’s dense points observed relative to the total dense points of that primitive.
-For instance proposals, we utilize Sparse Non-Maximum Suppression(Sparse-NMS) on the proposals generated from all windows, which increases processing efficiency (detailed analysis is provided in the supplementary material).
-A potential instance may be partially or fully observed across multiple windows.
-By ensuring a sufficiently small sliding step size, each instance is fully captured within at least one window.
-Incomplete proposals from other windows are subsequently filtered out during the NMS process.
-3.4 Automated 3D Interior Reconstruction
-CADSpotting delivers precise panoptic segmentation results, serving as the basis for our parametric 3D interior modeling algorithm.
-Utilizing detailed instance information and primitive positions from segmentation, we apply instance-level processing techniques to compute spatial parameters for key architectural elements, specifically walls, doors, and windows.
-For doors, we determine their positions, orientations, and pivot points by analyzing the relationships between arcs and lines within each segmented instance.
-Window positions, on the other hand, are directly extracted from their corresponding instances.
-Complex wall structures require a specialized approach: we first traverse and merge endpoints of adjacent lines within wall instances, forming closed polygons.
-These polygons are then rasterized into binary images, from which final wall contours are identified using the method from [32].
-With these spatial parameters, we reconstruct a complete 3D interior model using a parametric reconstruction pipeline in Blender.
-The necessary 3D primitives for walls, doors, and windows are manually prepared in Blender, providing reusable templates for streamlined parametric interior modeling.
-This approach significantly enhances efficiency, automating the conversion of raw CAD data into structurally accurate 3D interior models within minutes through procedural generation.
-An example result from our automated 3D interior reconstruction method is illustrated in Fig.
-4, with additional details provided in the supplementary material.
+The mask loss combines binary cross-entropy and Dice loss (Milletari, Navab, and Ahmadi 2016) , jointly measuring the alignment between predicted instance masks and ground truth.
+Sliding Window Aggregation
+To address the challenge of panoptic symbol segmentation at large scales, a key issue is generalizing instance segmentation across drawings of varying sizes.
+During inference, larger inputs require significantly more instance proposals than those seen during training.
+However, Mask Transformer-based methods, including ours, are constrained by a fixed number of instance queries.
+Fortunately, CAD floorplans offer an advantage: instances of the same category typically maintain consistent sizes across different scenes under uniform scaling.
+This allows us to assume that instance sizes at inference do not exceed those in training data.
+Based on this assumption, we adopt a fixed-size sliding window strategy, using windows aligned with training dimensions to process large input drawings.
+For each window, we collect all primitives it covers or intersects and apply the decoder to generate local semantic and instance proposals.
+Aggregated results across all windows are then merged using distinct strategies for semantic and instance outputs.
+For semantic proposals, we apply a weighted voting scheme to assign each primitive a class label.
+When a primitive appears in multiple windows, we prioritize proposals from windows that observe a larger portion of it.
+The voting weight is proportional to the number of observed dense points relative to the total points of the primitive.
+For instance proposals, we use Sparse Non-Maximum Suppression (Sparse-NMS) to merge results from overlapping windows efficiently (see supplementary material).
+By using a sufficiently small step size, each instance is fully captured in at least one window.
+Incomplete detections from other windows are eliminated during the NMS process.
+Automated 3D Interior Reconstruction
+CADSpotting produces accurate panoptic segmentation results that serve as the foundation for our parametric 3D interior modeling pipeline.
+Using instance-level segmentation and primitive positions, we extract spatial parameters for key architectural elements, including walls, doors, and windows.
+For doors, we compute positions, orientations, and pivot points by analyzing geometric relationships between arcs and lines within each segmented instance.
+Window positions are directly derived from their corresponding instances.
+Wall modeling involves a more specialized process: we merge endpoints of adjacent lines within wall instances to form closed polygons, which are rasterized into binary masks.
+Final wall contours are then extracted using the method from (Suzuki and be 1985).
+These spatial parameters are fed into a parametric reconstruction pipeline implemented in Blender.
+Predefined 3D assets for walls, doors, and windows are used as reusable templates, enabling efficient and automated generation of structurally accurate 3D interiors.
+This approach converts raw CAD inputs into detailed 3D models within minutes.
+An example result is shown in Fig.
+4, with further details provided in the supplementary material.
 | Building | Area() | #Primitive | #Instance |
@@
 | Hotel 1 | 1,193 | 26,709 | 335 |
-4 Experiments
-4.1 The Proposed LS-CAD Dataset
-We introduce LS-CAD, a new large-scale CAD dataset consisting of 50 floorplans from extensive buildings, such as campuses and office buildings.
-Each CAD drawing in LS-CAD covers an area of at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over ten thousand.
-Table.1 showcases 3 large-scale CAD drawings from our dataset, providing detailed information such as area, number of primitives and instance count, which underscore the large-scale nature of LS-CAD.
-Additionally, the floorplans in LS-CAD include a wide variety of complex symbols, such as doors, windows, walls, elevators and parking spaces.
-We provide fine-grained annotations consistent with the standards of the FloorPlanCAD [5] dataset.
-A portion of LS-CAD will be publicly available under a license waiver to support further research in panoptic symbol spotting for large-scale CAD drawings.
-4.2 Experimental Settings
-Implementation Details.
+Experiments
+The Proposed LS-CAD Dataset
+We introduce LS-CAD, a new large-scale CAD dataset comprising 45 floorplans from expansive buildings, such as campuses and office complexes.
+Each drawing covers at least 1,000 square meters, with the number of primitives ranging from approximately 2,900 to over 10,000.
+Table.
+1 presents three representative examples, detailing area, primitive count, and instance count to highlight the large-scale nature of LS-CAD.
+The dataset features a rich variety of complex symbols, including doors, windows, walls, elevators, and parking spaces.
+All floorplans are annotated with fine-grained labels consistent with the FloorPlanCAD (Fan et al.
+2021) standard.
+The LS-CAD dataset will be publicly released to support future research in panoptic symbol spotting for large-scale CAD drawings.
+Experimental Settings
+Implementation Details
 The FloorPlanCAD dataset comprises approximately 15,000 drawings, annotated across 35 symbol categories, including 30 things and 5 stuff classes.
@@
 For dense point sampling, we set a fixed distance .
-We use the AdamW [19] optimizer with a learning rate of and a weight decay of 0.05.
-Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes and color normalization to .
+We use the AdamW (Loshchilov 2017) optimizer with a learning rate of and a weight decay of 0.05.
+Data augmentation techniques include random horizontal flipping with a probability of , global rotation and scaling transformations, translation along the x and y axes.
 We set loss weight as .
@@
 We use as the step size of SWA on LS-CAD dataset.
-Metrics.
-Following the approach of [6], we assess the performance of our model using multiple metrics.
-For semantic symbol spotting, we use F1 score, weighted F1 score (wF1), and mean Intersection over Union (mIoU).
+Metrics
+Following the approach of (Fan et al.
+2022), we assess the performance of our model using multiple metrics.
+For semantic symbol spotting, we use F1 score and weighted F1 score (wF1).
 For instance symbol spotting, we employ AP50, AP75, and mean Average Precision (mAP).
 For panoptic symbol spotting, we utilize Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ).
-PQ is a comprehensive metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks.
-Further details on metric formulation are available in [6].
-4.3 Quantitative Evaluation
-| Methods | AP50 | AP75 | mAP |
-| DINO [38] | 64.0 | 54.9 | 47.5 |
-| SymPoint [16] | 66.3 | 55.7 | 52.8 |
-| CADSpotting (ours) | 72.2 | 69.1 | 69.0 |
+PQ is a metric combining SQ and RQ to evaluate model performance in semantic and instance segmentation tasks.
+Further details on metric formulation are available in (Fan et al.
+2022).
+Quantitative Evaluation
+| Method | AP50 | AP75 | mAP |
+| DINO (Zhang et al. 2022) | 64.0 | 54.9 | 47.5 |
+| SymPoint (LIU et al. 2024) | 66.3 | 55.7 | 52.8 |
+| CADSpotting (ours) | 70.6 | 66.6 | 66.3 |
 | Method | Total | Thing | Stuff | ||||||
 | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
-| PanCADNet [5] | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
-| CADTransformer [6] | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
-| GAT-CADNet [40] | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
-| SymPoint [16] | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
-| CADSpotting (Ours) | 88.9 | 95.6 | 93.0 | 89.7 | 96.2 | 93.2 | 80.6 | 89.7 | 89.8 |
-To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD [5] and LS-CAD datasets.
+| PanCADNet (Fan et al. 2021) | 59.5 | 82.6 | 66.9 | 65.6 | 86.1 | 76.1 | 58.7 | 81.3 | 72.2 |
+| CADTransformer (Fan et al. 2022) | 68.9 | 88.3 | 73.3 | 78.5 | 94.0 | 83.5 | 58.6 | 81.9 | 71.5 |
+| GAT-CADNet (Zheng et al. 2022) | 73.7 | 91.4 | 80.7 | - | - | - | - | - | - |
+| SymPoint (LIU et al. 2024) | 83.3 | 91.4 | 91.1 | 84.1 | 94.7 | 88.8 | 48.2 | 69.5 | 69.4 |
+| CADSpotting (Ours) | 87.4 | 93.5 | 93.4 | 88.3 | 94.2 | 93.7 | 71.5 | 82.9 | 86.3 |
+To demonstrate the effectiveness of our approach, we compare it against SOTA methods on the FloorPlanCAD (Fan et al.
+2021) and LS-CAD datasets.
 A detailed comparison of semantic, instance, and panoptic symbol spotting tasks is provided in the following paragraphs.
-Semantic Symbol Spotting.
-As shown in Table.2, we compare our method with PanCADNet [5], CADTransformer [6], GAT-CADNet [40], and SymPoint [16].
-Our CADSpotting approach achieves the highest performance in both the F1 and wF1 metrics, outperforming SymPoint by 13.6% in mIoU.
-Instance Symbol Spotting.
-We also evaluate our method against DINO [38] and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint.
-As shown in Table.3, our method achieves superior results in the AP50, AP75 and mAP metrics, with improvements of 5.9% AP50, 8.4% AP75 and 8.9% mAP over SymPoint.
-Panoptic Symbol Spotting.
-We then compare our method with the same methods used in the semantic symbol spotting comparison.
+Semantic Symbol Spotting As shown in Table.2, we compare our method with PanCADNet (Fan et al.
+2021), CADTransformer (Fan et al.
+2022), GAT-CADNet (Zheng et al.
+2022), and SymPoint (LIU et al.
+2024).
+Our CADSpotting achieves the highest performance in both the F1 and wF1 metrics.
+Instance Symbol Spotting We also evaluate our method against DINO (Zhang et al.
+2022) and SymPoint for instance symbol spotting, using the bounding box maximization of predicted mask to compute the AP metric, consistent with SymPoint.
+As shown in Table.3, our method achieves superior results , with improvements of 4.3 p.p.
+AP50, 10.9 p.p.
+AP75 and 13.5 p.p.
+mAP over SymPoint.
+Panoptic Symbol Spotting We compare our method with the same methods used in the semantic symbol spotting comparison.
 The results, presented in Table.4, show that our approach surpasses SymPoint and other prior methods across all metrics, including Total/Thing/Stuff PQ, SQ and RQ.
 Additionally, our approach offers superior generalization across diverse CAD primitives, thanks to its dense point sampling based feature representation.
-We also compare the effectiveness of Sliding Window Aggregation (SWA) and Block Partitioning (BP) methods when applied to the CADSpotting model on the LS-CAD dataset.
-We apply BP method using the same size as drawings the FloorPlanCAD dataset without overlapping between blocks, to divide large-scale CAD drawings into smaller blocks.
-When employing the BP method, we ensure that each primitive is uniquely represented by retaining only those whose starting points fall within the block.
-The results in Table 5 indicate that, compared to the BP method, SWA achieves the highest performance on PQ, with an absolute improvement of 10.4%, although there is a slight decrease of 1.0% on SQ.
-The result shows that our CADSpotting method with the SWA technique can accurately recognize instances of CAD symbols in large-scale CAD drawings.
-4.4 Qualitative Comparison
+| Method | Panoptic | Semantic | Instance | ||
+| PQ | SQ | RQ | F1 | mAP | |
+| SPv2 | 90.1 | 96.3 | 93.6 | 89.5 | 60.1 |
+| Ours | 87.4 | 93.5 | 93.4 | 92.8 | 66.3 |
+| Method | Panoptic | Semantic | Instance | ||
+| PQ | SQ | RQ | F1 | mAP | |
+| SPv2 + BP | 26.7 | 82.0 | 32.6 | 55.2 | 15.1 |
+| Ours + BP | 58.8 | 83.7 | 70.2 | 88.7 | 48.3 |
+| Ours + SWA | 75.5 | 80.9 | 93.3 | 93.5 | 57.5 |
+We then compare our method with the concurrent work SymPoint-V2 (SPv2) (Liu et al.
+2024), which incorporates primitive types into feature vectors.
+Our first comparison is conducted on the FloorPlanCAD dataset.
+As shown in Table.5, although SPv2 achieves up to 2.7 percentage points higher PQ in panoptic segmentation, our method delivers a 3.3-point gain in F1 score and a 6.2-point improvement in mAP.
+While SPv2 performs well on FloorPlanCAD due to its reliance on four fixed primitive types, this limits its generalization to more diverse CAD symbol sets.
+In contrast, our method relies solely on coordinate attributes, enabling more flexible and robust feature learning.
+Furthermore, SPv2 depends on CAD layer information, which is often inconsistent or unavailable in real-world scenarios.
+Table.
+6 further supports this analysis.
+On the LS-CAD dataset, where layer annotations are unavailable, our method significantly outperforms SPv2 across all metrics, demonstrating stronger robustness in practical applications.
+We also evaluate the effectiveness of Sliding Window Aggregation (SWA) versus Block Partitioning (BP) in CADSpotting on the LS-CAD dataset.
+In the BP setting, we divide large-scale CAD drawings into non-overlapping blocks matching the dimensions used in FloorPlanCAD.
+To ensure unique primitive representation, only those with starting points inside a block are retained.
+As shown in Table 6, SWA substantially outperforms BP in panoptic quality (PQ), achieving a 16.7-point improvement, with only a modest 2.8-point reduction in segmentation quality (SQ).
+These results demonstrate that CADSpotting, when combined with SWA, achieves accurate and scalable symbol detection in large-scale CAD drawings.
+Qualitative Comparison
 We perform qualitative comparisons with SymPoint on the FloorPlanCAD dataset.
@@
 These results demonstrate that our CADSpotting enhanced by SWA technique, accurately identifies semantic and instance information in large-scale CAD drawings.
-| Method | PQ | SQ | RQ |
-| Ours + BP | 42.8 | 94.8 | 45.2 |
-| Ours + SWA | 53.2 | 93.8 | 56.6 |
 To evaluate model generalizability and demonstrate the utility of LS-CAD dataset, we also train models on a combined dataset consisting of FloorPlanCAD and LS-CAD.
 Comprehensive experimental results and analyses are provided in the supplementary material.
-| Methods | Backbone | PQ | SQ | RQ |
+| Method | Backbone | PQ | SQ | RQ |
 | Handcrafted Features | PTv1 | 78.4 | 88.5 | 88.6 |
 | Handcrafted Features | PTv3 | 80.2 | 90.2 | 89.0 |
-| Ours | PTv3 | 88.9 | 95.6 | 93.0 |
-| Method | PQ | SQ | RQ |
-| w/o Color Feature | 85.8 | 92.5 | 92.7 |
-| Color Feature | 88.9 | 95.6 | 93.0 |
-4.5 Ablation Studies
-In this section, we present ablation studies examining the effects of feature learning and RGB information.
-Additional ablation studies, which explore various pooling strategies and evaluate performance across all classes, are provided in the supplementary material.
-Feature Learning.
-We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive.
-SymPoint employs Point Transformer(PTv1) [39] as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance.
-To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 [35], maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting.
-As shown in Table.6, with the same decoder configuration, simply switching to PTv3 results in a 1.8% improvement in PQ.
-When we apply our proposed primitive feature learning method, performance further increases by 8.7% in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
-RGB Color Information.
-We compare the performance of CADSpotting through the inclusion versus exclusion of color information in the FloorPlanCAD dataset.
-Table.7 shows that adding RGB information improves the PQ metric by 3.1%.
-5 Conclusion
+| Ours | PTv3 | 87.4 | 93.5 | 93.4 |
+Ablation Studies
+We further present ablation studies on feature learning.
+Additional ablation studies, which explore various pooling strategies and the use of color information, and evaluate performance across all classes, are provided in the supplementary material.
+Feature Learning We compare our dense point sampling based primitive feature learning method with the approach in SymPoint, which uses handcrafted features to represent each individual primitive.
+SymPoint employs Point Transformer(PTv1) (Zhao et al.
+2021) as its backbone network and incorporates hierarchical multi-resolution primitive features to leverage intermediate features, enhancing the decoder’s performance.
+To evaluate the impact of different point cloud analysis methods, we first replace SymPoint’s backbone with PTv3 (Wu et al.
+2024), maintaining the same decoder architecture as ours to ensure a fair comparison of feature learning techniques in panoptic symbol spotting.
+As shown in Table.7, with the same decoder configuration, simply switching to PTv3 results in a 1.8 p.p.
+improvement in PQ.
+When we apply our proposed primitive feature learning method, performance further increases 7.2 p.p.
+in PQ, demonstrating the significant performance gains achieved by our dense point sampling based feature learning approach.
+Conclusion
 We have introduced CADSpotting, a novel method for panoptic symbol spotting in large-scale CAD drawings.
 Our approach uses dense point sampling within a unified point cloud framework to represent CAD primitives, combined with a Sliding Window Aggregation technique that integrates weighted voting and Non-Maximum Suppression for precise segmentation.
-We also present LS-CAD, a dataset of 50 annotated floorplans from diverse building types.
+We also present LS-CAD, a dataset of 45 annotated floorplans from diverse building types.
 Experiments on LS-CAD and FloorPlanCAD datasets have confirmed the effectiveness and scalability of our method, with successful applications in automated 3D reconstruction demonstrating its practical value.
@@
 | PQ | SQ | RQ | PQ | SQ | RQ | ||
-| SymPoint [16] | FloorPlanCAD | 83.3 | 91.4 | 91.1 | 33.2 | 83.0 | 39.9 |
+| SymPoint (LIU et al. 2024) | FloorPlanCAD | 83.3 | 91.4 | 91.1 | 33.2 | 83.0 | 39.9 |
 | SymPoint | Cross-dataset | 79.0 | 89.2 | 88.6 | 57.0 | 84.0 | 67.8 |
-| CADSpotting | FloorPlanCAD | 88.9 | 95.6 | 93.0 | 42.8 | 94.8 | 45.2 |
-| CADSpotting | Cross-dataset | 87.3 | 95.4 | 91.4 | 69.9 | 91.6 | 76.3 |
-| CADSpotting(SWA) | FloorPlanCAD | / | 53.2 | 93.8 | 56.6 | ||
-| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 84.9 | 88.9 | ||
-6 More Details on the LS-CAD Dataset
+| CADSpotting | FloorPlanCAD | 87.4 | 93.5 | 93.4 | 52.9 | 90.4 | 58.5 |
+| CADSpotting | Cross-dataset | 84.0 | 91.3 | 91.9 | 58.8 | 83.7 | 70.2 |
+| CADSpotting(SWA) | FloorPlanCAD | / | 60.1 | 90.9 | 66.1 | ||
+| CADSpotting(SWA) | Cross-dataset | / | 75.5 | 80.9 | 93.3 | ||
+More Details on the LS-CAD Dataset
 In Fig.8, we present visualizations of several samples from the LS-CAD dataset to highlight the diversity of the proposed dataset.
@@
 Notably, CADSpotting+SWA method reaches the best performance of 75.5 on the PQ metric.
-After cross-dataset joint training, the PQ score for CADSpotting decreases by only 1.2% on FloorPlanCAD, whereas it drops by 4.3% on SymPoint.
+After cross-dataset joint training, the PQ score for CADSpotting decreases by only 3.4 p.p.
+on FloorPlanCAD, whereas it drops by 4.3 p.p.
+on SymPoint.
 This indicates that our model exhibits stronger generalization capabilities.
-7 Quantitative Evaluation
+Quantitative Evaluation
 Ablation Study on Feature Learning.
@@
 We further evaluate the necessity and impact of different pooling methods on our approach, using a baseline configuration without a pooling layer.
-As shown in Table.9, max pooling improves performance by 11.5%, average pooling by 13.0%, and mixed pooling by 14.3% in PQ.
+As shown in Table.9, max pooling improves performance by 5.0 p.p., average pooling by 5.1 p.p., and mixed pooling by 5.5.
+p.p.
+in PQ.
 Based on these results, we select mixed pooling for our CADSpotting approach.
-8 Qualitative Evaluation
+Ablation Study on Prior Color Information.
+To evaluate the necessity of color prior, we incorporate RGB channels as an optional input.
+As shown in Table 10, adding color improves PQ by 1.5 p.p.
+and SQ by 2.1 p.p., while slightly decreasing RQ by 0.4 p.p.
+More importantly, the geometry-based representation (w/o Color) achieves strong baseline performance, demonstrating that our model maintains high recognition quality even without color information.
+Qualitative Evaluation
 We provide comparisons of prediction results on different datasets.
@@
 For LS-CAD, we present the results of semantic and panoptic predictions, highlighting the performance of the method in practical scenarios, as shown in Fig.11 and Fig.12.
-9 Sliding Window Implementation Details
+Sliding Window Implementation Details
 The Algorithm.1 details the implementation process of our SWA.
-In the matrix-NMS [33] process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix.
+In the matrix-NMS (Wang et al.
+2020) process for sliding window aggregation, computing the Intersection over Union (IoU) between all pairs of instance proposal masks is essential for constructing the IoU matrix.
 However, the number of instance proposals grows substantially with the size of the input drawing.
@@
 By leveraging this sparsity, we reimplemented the IoU computation using sparse matrix routines, significantly reducing computational cost and memory usage without sacrificing accuracy.
-10 Automated 3D Interior Reconstruction
+Automated 3D Interior Reconstruction
 Based on the spotting results generated by CADSpotting, we present more 3D reconstruction renderings in Fig.14.
@@
 | Primitive Pooling Type | PQ | SQ | RQ |
-| w/o Pooling | 74.6 | 89.3 | 83.6 |
-| Max | 86.1 | 94.1 | 91.5 |
-| Average | 87.6 | 94.5 | 92.7 |
-| Mixed (Max+Average) | 88.9 | 95.6 | 93.0 |
+| w/o Pooling | 81.9 | 89.4 | 91.6 |
+| Max | 86.9 | 93.3 | 93.2 |
+| Average | 87.0 | 93.1 | 93.4 |
+| Mixed (Max+Average) | 87.4 | 93.5 | 93.4 |
+| Method | PQ | SQ | RQ |
+| w/ color | 87.4 | 93.5 | 93.3 |
+| w/o color | 88.9 | 95.6 | 93.0 |
 | class | A | B | C | D | ||||||||
 | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | PQ | SQ | RQ | |
-| single door | 91.59 | 97.08 | 94.34 | 75.67 | 90.68 | 83.44 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
-| double door | 93.51 | 96.93 | 96.47 | 75.14 | 88.76 | 84.65 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
-| sliding door | 96.54 | 98.78 | 97.73 | 83.55 | 50.31 | 83.55 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
-| folding door | 57.67 | 87.88 | 65.62 | 39.66 | 95.83 | 41.38 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
+| single door | 91.90 | 96.33 | 95.40 | 84.11 | 91.34 | 92.08 | 87.23 | 92.68 | 94.12 | 87.46 | 92.85 | 94.20 |
+| double door | 94.10 | 96.46 | 97.55 | 87.23 | 91.70 | 95.13 | 87.30 | 91.85 | 95.04 | 88.27 | 92.73 | 95.18 |
+| sliding door | 96.50 | 98.02 | 98.46 | 92.88 | 95.04 | 97.73 | 94.51 | 96.15 | 98.29 | 91.32 | 95.65 | 95.46 |
+| folding door | 77.06 | 89.23 | 86.36 | 73.03 | 92.36 | 79.07 | 70.88 | 83.29 | 85.11 | 52.03 | 82.64 | 62.96 |
 | revolving door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
 | rolling door | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
-| window | 89.60 | 97.24 | 92.14 | 75.49 | 91.87 | 82.17 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
-| bay window | 48.15 | 95.26 | 50.55 | 41.92 | 88.51 | 47.37 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
-| blind window | 90.57 | 98.17 | 92.25 | 77.71 | 96.14 | 80.83 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
-| opening symbol | 46.64 | 76.82 | 60.71 | 37.97 | 72.60 | 52.30 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
-| sofa | 88.53 | 96.29 | 91.93 | 71.82 | 82.05 | 87.53 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
-| bed | 89.79 | 92.47 | 97.11 | 78.30 | 83.76 | 93.47 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
-| chair | 83.58 | 94.04 | 88.88 | 70.05 | 85.54 | 81.89 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
-| table | 75.39 | 88.46 | 85.22 | 53.18 | 78.36 | 67.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
-| TV cabinet | 95.23 | 97.57 | 97.60 | 82.85 | 86.79 | 95.46 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
-| Wardrobe | 93.76 | 97.23 | 96.43 | 84.53 | 89.58 | 94.36 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
-| cabinet | 81.40 | 91.16 | 89.28 | 65.83 | 81.45 | 80.82 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
-| gas stove | 97.34 | 99.17 | 98.16 | 90.08 | 96.15 | 93.69 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
-| sink | 87.59 | 95.21 | 92.00 | 76.13 | 91.10 | 83.57 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
-| refrigerator | 91.96 | 96.24 | 95.55 | 71.86 | 81.87 | 87.78 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
-| air conditioner | 86.53 | 99.02 | 87.39 | 78.66 | 96.36 | 81.64 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
-| bath | 79.30 | 93.31 | 84.98 | 63.84 | 83.21 | 76.73 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
-| bath tub | 85.00 | 91.92 | 92.47 | 70.96 | 82.90 | 85.61 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
-| washing machine | 89.17 | 97.74 | 91.23 | 78.29 | 90.75 | 86.27 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
-| squat toilet | 95.35 | 98.25 | 97.05 | 78.70 | 94.34 | 83.42 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
-| urinal | 94.32 | 98.76 | 95.51 | 80.33 | 92.08 | 87.24 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
-| toilet | 95.21 | 98.08 | 97.07 | 81.39 | 91.92 | 88.55 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
-| stairs | 86.47 | 94.28 | 91.72 | 74.68 | 89.68 | 83.27 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
-| elevator | 96.15 | 97.53 | 98.58 | 79.08 | 94.76 | 83.44 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
-| escalator | 73.73 | 88.37 | 83.44 | 67.98 | 79.15 | 85.88 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
-| row chairs | 85.16 | 94.94 | 89.71 | 82.45 | 92.43 | 89.21 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
-| parking spot | 89.68 | 95.20 | 94.20 | 84.34 | 92.67 | 91.01 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
-| wall | 82.37 | 88.83 | 92.73 | 75.08 | 81.61 | 66.81 | 81.86 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
-| curtain wall | 64.80 | 89.88 | 72.09 | 64.17 | 84.45 | 75.98 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
-| railing | 75.67 | 92.22 | 82.06 | 62.87 | 85.69 | 73.37 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
-| total | 88.91 | 95.65 | 92.96 | 74.59 | 89.26 | 83.56 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |
+| window | 85.41 | 91.06 | 93.80 | 77.71 | 84.43 | 92.04 | 77.13 | 85.24 | 90.49 | 74.53 | 85.06 | 87.62 |
+| bay window | 51.18 | 90.19 | 56.74 | 43.23 | 91.50 | 47.25 | 22.25 | 77.38 | 28.75 | 12.00 | 68.83 | 17.44 |
+| blind window | 86.00 | 92.25 | 93.22 | 77.99 | 86.97 | 89.68 | 79.56 | 84.85 | 93.76 | 77.28 | 83.91 | 92.10 |
+| opening symbol | 52.67 | 77.06 | 68.35 | 34.51 | 74.91 | 46.07 | 36.47 | 79.11 | 46.10 | 20.10 | 72.22 | 27.84 |
+| sofa | 89.68 | 96.78 | 92.67 | 85.50 | 92.42 | 92.51 | 76.98 | 90.28 | 85.28 | 73.74 | 87.40 | 84.37 |
+| bed | 82.12 | 88.75 | 92.54 | 79.51 | 86.32 | 92.11 | 76.87 | 86.26 | 89.11 | 76.51 | 84.44 | 90.61 |
+| chair | 84.72 | 93.66 | 90.46 | 86.51 | 94.63 | 91.42 | 84.42 | 92.26 | 91.51 | 80.85 | 91.71 | 88.16 |
+| table | 77.42 | 89.58 | 86.42 | 73.30 | 87.39 | 83.87 | 68.03 | 86.46 | 78.68 | 61.99 | 86.43 | 71.72 |
+| TV cabinet | 94.56 | 97.08 | 97.40 | 83.23 | 85.80 | 97.01 | 89.99 | 93.41 | 96.33 | 79.94 | 85.50 | 93.49 |
+| Wardrobe | 91.56 | 94.54 | 96.85 | 88.97 | 92.63 | 96.05 | 85.37 | 88.77 | 96.17 | 82.50 | 85.94 | 95.99 |
+| cabinet | 80.10 | 89.04 | 89.95 | 73.11 | 84.25 | 86.78 | 71.37 | 82.66 | 86.33 | 69.15 | 83.42 | 82.90 |
+| gas stove | 97.89 | 99.21 | 98.67 | 96.76 | 98.39 | 98.35 | 97.10 | 97.89 | 99.19 | 96.66 | 97.61 | 99.03 |
+| sink | 87.08 | 93.71 | 92.93 | 84.87 | 91.49 | 92.77 | 84.88 | 90.98 | 93.30 | 83.42 | 90.03 | 92.65 |
+| refrigerator | 91.05 | 95.32 | 95.52 | 80.35 | 84.60 | 94.97 | 81.70 | 87.44 | 93.44 | 80.28 | 85.58 | 93.80 |
+| air conditioner | 87.43 | 98.17 | 89.06 | 82.21 | 95.67 | 85.93 | 80.59 | 92.04 | 87.56 | 75.45 | 91.70 | 82.27 |
+| bath | 75.50 | 88.30 | 85.50 | 69.96 | 81.78 | 85.54 | 66.62 | 79.71 | 83.58 | 62.06 | 77.36 | 80.22 |
+| bath tub | 83.67 | 88.14 | 94.92 | 76.50 | 81.37 | 94.01 | 69.94 | 80.14 | 87.27 | 64.82 | 78.13 | 82.96 |
+| washing machine | 88.26 | 95.38 | 92.54 | 82.34 | 88.59 | 92.95 | 86.40 | 91.21 | 94.73 | 77.97 | 85.26 | 91.45 |
+| squat toilet | 92.85 | 96.01 | 96.71 | 92.41 | 96.23 | 96.03 | 92.58 | 94.79 | 97.67 | 89.99 | 92.68 | 97.09 |
+| urinal | 94.46 | 97.76 | 96.62 | 91.24 | 94.17 | 96.90 | 91.80 | 94.84 | 96.79 | 91.31 | 94.79 | 96.32 |
+| toilet | 95.83 | 97.70 | 98.09 | 91.22 | 92.96 | 98.13 | 90.81 | 93.01 | 97.63 | 90.48 | 93.53 | 96.74 |
+| stairs | 84.41 | 92.37 | 91.38 | 80.19 | 88.51 | 90.60 | 68.56 | 82.02 | 83.59 | 65.50 | 79.52 | 82.36 |
+| elevator | 95.06 | 97.57 | 97.43 | 93.26 | 96.46 | 96.69 | 84.44 | 90.52 | 93.28 | 80.64 | 88.59 | 91.03 |
+| escalator | 74.96 | 87.62 | 85.55 | 70.43 | 85.84 | 82.05 | 29.80 | 72.72 | 40.98 | 44.64 | 73.70 | 60.56 |
+| row chairs | 85.73 | 94.58 | 90.65 | 84.15 | 93.96 | 89.55 | 85.69 | 94.53 | 90.65 | 84.45 | 93.99 | 89.86 |
+| parking spot | 88.54 | 93.15 | 95.05 | 84.48 | 90.24 | 93.62 | 72.22 | 80.84 | 89.34 | 72.64 | 82.77 | 87.76 |
+| wall | 73.21 | 81.42 | 89.91 | 62.55 | 73.32 | 85.31 | 44.88 | 65.29 | 68.74 | 44.48 | 65.44 | 67.97 |
+| curtain wall | 58.66 | 82.61 | 71.01 | 50.49 | 78.00 | 64.73 | 60.25 | 87.36 | 73.90 | 37.62 | 73.21 | 51.39 |
+| railing | 67.25 | 88.38 | 76.10 | 59.50 | 83.88 | 70.94 | 37.82 | 76.42 | 49.49 | 28.53 | 70.73 | 40.34 |
+| total | 87.36 | 93.40 | 93.54 | 81.89 | 89.42 | 91.58 | 80.24 | 90.17 | 88.98 | 78.36 | 88.59 | 88.46 |
 A: Dense point sampling + PTv3 + Pooling(Ours). B: Dense point sampling + PTv3. C: Handcrafted features + PTv3. D: Handcrafted features + PTv1